Solve the system by addition. 4x – 6y = –1 8x + 3y = 3

Answer by checkley71(8403)

4X-6Y=-1 NOW MULTIPLY THE OTHER EQUATION BY 2 & ADD THEM.
16X+6Y=6
------------------------
20X=5
X=5/20
X=1/4 ANSWER.
4*1/4-6Y=-1
1-6Y=-1
-6Y=-1-1
-6Y=-2
Y=-2/-6
Y=1/3 ANSWER.
PROOF
8*1/4+3*1/3=3
8/4+3/3=3
2+1=3
3=3